OpenStudy (anonymous):

Faith mixes 14 liters of 12% acid solution with a 30% acid solution which results in an 18% acid solution. Find the number of liters of 30% solution in the new mixture.

hero (hero):

x = amount of second solution y = total amount of mixed solution 14 + x = y 14(.12) + x(.30) = y(.18)

hero (hero):

Solve the system

OpenStudy (anonymous):

.30x - 0.18y = -1.68 x - y = -14 -0.18x + 0.18y = 2.52 add 1st and 3rd eqns 0.12x = 2.52-1.68 = 0.84 x = 7 14+7 = y = 21 check 21 (.18) = 3.78 14(.12) + 7(.30) = 3.78 right